In some 4.0 grading systems, a student's grade point average (GPA) is calculated by assigning letter grades the following numerical values. 01 points AufExo4 13.1.019. 0100 s In some 4.0 greding systems, a students grade point average (GPA) is calculated by assign
Answer #1

Let the first course unit be x1 , second be x2 and so on..

Therefore,

(X1 + x2 + ….. + x46 )/46 = 3.19

(X1 + x2 + ….. + x46 ) = 3.19*46 = 146.74

Also,

(x47 + x48 + ……+x58)/ 12 = 3.8

(x47 + x48 + ……+x58) = 45.6

Therefore, Tessa’s cumulative GPA fro all 4 semesters will be the sum of all course units divided by 58(46+ 12)

Therefore,

(X1 + x2 + ….. + x46 + x47 + x48 + ……+x58)/ 58 = (146.74 + 45.6)/ 58 = 192.34/58 = 3.316

Hence her GPA for all 4 semesters- 3.32